Refer to Dino's information below and calculate the following ratios: Working capital ratio, quick ratio/acid test, earnings per share, price-earnings ratio, debt-equity ratio, and return on equity. Clearly provide each formulae, numbers and work associated, along with the answers.

Dinos After School Balance Sheet Dec 31 20XX Current Assets: Cash Accts Receivable invento Supplies Marketable securities Total Current assets Liabilities: $2,500 Accts payable $9,000 Wages payable $2,000 Taxes payable $6,000 $1,000 $2,000 $500 Unearned revenue $4,000 $10,000 $24,000 Total current liab $13,000 Long term liabilities Mortgage payable$75,000 Bank note payable$20,000 Long term assets: Building Equipment less: accumulated depreciation $150,000 Total liabilities $70,000 ($60,000) $108,000 Owners equity Common stock $10,000 $65,000 Retained Earnings Total equity Total liabilities plus Total Long term assets $160,000 $76,000 Total assets $184,000 Additional information: #Of shares outstanding 10,000 share price Net income Dividends were.50 per share owners equity $184,000 $20 $12,000

For each ratio you calculated, provide an analytic statement in which you comment on how the ratio can affect specific decisions that need to be made within the organization.

Notes:

  • Each statement must be at least one full paragraph, explaining the ratio results in isolation along the scale, then application to the specific business.
  • Specifically include throughout your analyses, decisions regarding inventory management, depreciation of company assets, employee wages, government taxes, and investment prospects.
  • 75-150 words for each ratio.
Answer #1

Working capital ratio = current assets / current liabilities

= 24000/13000 = 1.85.

Current ratios show for $1 current liability to pay, how much current assets are available. In the present case for each $1 current liability to be paid we have $1.85 worth of current assets. There is no standard regarding current ratio. However anything between 1.2 to 2 is considered to be ideal. Thus, company is having ample current assets available with it for current liability . And thus short term liquidity of the company is strong and thus, company would try to maintain it's policies towards current assets and liabilities to be more and less same subject to the results of quick ratio.

2 . Quick ratio / acid test= current assets - inventories / current liabilities

= 24000-2000/13000 = 1.69.

Quick ratio shows how much current assets except inventory is available with the company to pay off its current liabilities. Inventories are excluded because the purpose here is to see how much current asset which is readily convertible into Cash is available with the company to pay off its current liabilities. In the present case we have around 1.69 of quick assets against $1 of current liability and thus it's quite strong enough position. Company would be happy to maintain it same on near future also.

3 . Earning per share = net income / outstanding number of shares

= 12000/10000 = 1.2

Earning per share goes to show that how much earning is attributable to each share of the company. In the given case it is around $1.2. we try to ensure that earning per share are high and accordingly make business decisions to ensure it remains high.

4. Price earning ratio = market price per share / earning per share

= 20/1.2 = 16.67

Price earning ratio shows for each $ earned what is it's impact on the market price of the share. In the present case it is 16.67 which means $ 1 earning changes market price by around 16.67. Decisions are made so that there will be maximum increase in market share as a result of earning per share.

5 . Debt-equity ratio = total debt / total stockholder equity

= 108000/76000 = 1.42.

Debt equity ratio shows the proportion of debt to the equity of the company. This ratios shows the Margin available for debtholders . A higher level of debt equity ratios is not considered very good as it means there is less Margin available with the debt holders to bear bear any default by the company. I'm the present case it is around 1.42 which is not very good. Companies generally ensure that it remains less than 1 as if it is high it also increase rate of interest on any incremental borrowing.

6 . Return on equity = net income / total equity x 100

= 12000/76000x100 = 15.79%.

It shows how much Return company is giving to its equityholders aa a compensation of investing their money in the Company. Greater the Return , greater the benefits for equity holder. Thus efforts are made to keep Return on equity always high.